# First steps towards Computational Polynomials in Lean

### v1

@@ II-A A preliminary choice
General-purpose representation in computer algebra systems are always sparse, otherwise one looks stupid if you can’t handle $x^{1000000000}+1$
@@ III Choices of Multivariate Polynomial Representations
Recursive. A typical representation for, say, $x^{3}-2x$ would be $(x,(3,1),(0,-2))$, i.e. a list starting with the variable, then ordered pairs as in “Sparse” above. In a typed language we might have a record type [variable,list]. Hence $z^{2}(y^{2}+2)+(3y+4)\in R[y][z]$ would be represented as
@@
A practical note. Experience in Axiom, as in [DGT91], shows that it may be useful to be able to talk about univariate polynomials in an unspecified variable, i.e. just a list of (exponent,coefficient) pairs with no variable specified.
@@ III-A Addition of Sparse polynomials
The notation xx@((i, a) :: x means “call it xx, but also deconstruct it into (i, a) as the head, and x as the tail.
Note that Lean requires a recursive definition to be proven to terminate. Any programmer would say that termination is obvious, as every recursive call is on less, but Lean doesn’t recognise this, as it says below.
@@
argument #5 was not used for structural recursion
@@
argument #6 was not used for structural recursion
@@
structural recursion cannot be used
The solution is
termination_by xx yy => xx.length + yy.length
@@ III-B Polynomial Multiplication
Naive polynomial multiplication (but see §IV) is relatively simple to implement given addition. Similarly the proofs that this algorithm gives well-formed results should follow from those results for addition. We have yet to see what the challenge is for correctness.
IV A new (to me) problem
@@ IV-A A specialist input mechanism?
Note that, at least in the worst case, the $\lambda_{i}$ may be much larger than $f$ or $f_{i}$. If the $\lambda_{i}$ are output in decreasing order, then the naive algorithm “read a term; add it to the polynomial” using the method of §III-A, will take $O((\#\lambda_{i})^{2})$ operations. A geobucket implementation (§II-B) would make this $O(\#\lambda_{i}\log(\#\lambda_{i}))$. However, an “intelligent” algorithm that knew it was reading in a specific order would be simply $O(\#\lambda_{i})$.
@@
We use a (balanced) binary heap to merge each (unevaluated) $f_{i}\times g$, so the sort key is $\deg(f_{i})+\deg(g)$, and the heap has #$\#f$ entries. After an $f_{i}\times g_{j}$ is extracted, we insert $f_{i}\times(g_{j+1}+\cdots)$. If the leading monomial of the resulting heap is the same as the just-extracted $f_{i}\times g_{j}$, we add the coefficients and continue. Since the size of the heap is #$\#f$, the cost of rebalancing after a single extraction is $\log\#f$, the number of extractions is #$\#f\#g$, and therefore the total cost is #$\#f\#g\log\#f$. We could therefore compute each $\lambda_{i}f_{i}$ this way, with $\lambda_{i}$ being $g_{i}$ as we might expect #$\#\lambda_{i}>\#f_{i}$ at least in the worst case.
IV-C New(?) idea
@@
A further improvement, rather than verifying $f=\sum^{N}f_{i}\lambda_{i}$, is to verify $0=(-1)f+\sum^{N}f_{i}\lambda_{i}$, so there is no output to construct.

### v2

@@ II-A A preliminary choice
General-purpose representation in computer algebra systems are always sparse, otherwise you look stupid if you can’t handle $x^{1000000000}+1$
@@ II-C An improvement
II-D Related work
The geobucket structure can be seen as deferring the expensive addition of differently-sized objects until the last possible moment. In this sense it has similarities to the Hierarchical Representation of [ZCJM06], as improved in [SLB24]. The principal difference is that in the hierarchical representation the user decides when to “unveil” (i.e. use the internal structure) an object, whereas in geobuckets the system automatically unveils the appropriately sized bucket, and actively updates the buckets. Hence the two systems have different utilities.
@@ III Choices of Multivariate Polynomial Representations
Recursive. A typical representation for, say, $x^{3}-2x$ would be $(x,(3,1),(1,-2))$, i.e. a list starting with the variable, then ordered pairs as in “Sparse” above. In a typed language we might have a record type [variable,list]. Hence $z^{2}(y^{2}+2)+(3y+4)\in R[y][z]$ would be represented as
@@
Experience in Axiom, as in [DGT91], shows that it may be useful to be able to talk about univariate polynomials in an unspecified variable, i.e. just a list of (exponent,coefficient) pairs with no variable specified.
@@ III-A Addition of Sparse polynomials
The notation xx@((i, a) :: x) means “call it xx, but also deconstruct it into (i, a) as the head, and x as the tail.
Note that Lean requires a recursive definition to be proven to terminate. Any programmer would say that termination is obvious, as every recursive call is on less (either less x or less y, or possibly both), but Lean doesn’t recognise this, as it says below (where we have used “R” for “recursion”).
@@
argument #5 was not used for structural R
@@
argument #6 was not used for structural R
@@
structural R cannot be used
[McK24] reports that an “obvious” translation into Agda has a similar issue with automatic proof of termination, and it needs a “non-obvious” translation to allow the automatic termination prover to conclude termination. Proofs of termination in Agda are discussed in [Agd24].
The author’s Lean solution is
termination_by xx yy =>
    xx.length + yy.length
@@ III-B Polynomial Multiplication
Naive polynomial multiplication (but see §IV-B) is relatively simple to implement given addition. Similarly the proofs that this algorithm gives well-formed results should follow from those results for addition. We have yet to see what the challenge is for correctness.
IV Polynomial Verification
@@ IV-A A specialist input mechanism?
Note that, at least in the worst case, the $\lambda_{i}$ may be much larger than $f$ or $f_{i}$. For example, in proving non-singularity of Weierstrass-form elliptic curves [Mac24, §3.2], the largest $f_{i}$ is $-y^{2}+2pxz+3qz^{2}i:=f_{3}$, but the largest $\lambda_{i}$ is
$\begin{array}[]{c}(256/3)p^{12}x^{2}+128qp^{11}xz+2304q^{2}p^{9}x^{2}+2592q^{3}p^{8}xz-\\ 64qp^{10}y^{2}+23328q^{4}p^{6}x^{2}+17496q^{5}p^{5}xz-1296q^{3}p^{7}y^{2}+\\ 104976q^{6}p^{3}x^{2}+39366q^{7}p^{2}xz-8748q^{5}p^{4}y^{2}+\\ 177147q^{8}x^{2}-19683q^{7}py^{2}=:\lambda_{1}\end{array}$
with 13 terms, and in fact $f_{1}\lambda_{1}$ has 26 terms.
If the $\lambda_{i}$ are output in decreasing order, then the naive algorithm “read a term; add it to the polynomial” using the method of §III-A, will take $O((\#\lambda_{i})^{2})$ operations. A geobucket implementation (§II-B) would make this $O(\#\lambda_{i}\log(\#\lambda_{i}))$. However, an “intelligent” algorithm that knew it was reading in a specific order would be simply $O(\#\lambda_{i})$.
@@
We use a (balanced) binary heap to merge each (unevaluated) $f_{i}\times g$, so the sort key is $\deg(f_{i})+\deg(g)$, and the heap has #$\#f$ entries. After an $f_{i}\times g_{j}$ is extracted, we insert $f_{i}\times(g_{j+1}+\cdots)$. If the leading monomial of the resulting heap is the same as the just-extracted $f_{i}\times g_{j}$, we add the coefficients and continue. Since the size of the heap is #$\#f$, the cost of rebalancing after a single extraction is $\log\#f$, the number of extractions is #$\#f\#g$, and therefore the total cost is #$\#f\#g\log\#f$.
We could therefore solve the problem from [Buz24b] by computing each $\lambda_{i}f_{i}$ this way, with $\lambda_{i}$ being $g_{i}$ as we might expect #$\#\lambda_{i}>\#f_{i}$ at least in the worst case. This would still requiring storing each $\lambda_{i}f_{i}$ (but if we computed $\sum\lambda_{i}f_{i}$ as we went along, we would store the partial sum and the single $\lambda_{i}f_{i}$ being added).
IV-C Verification without product construction
@@
A further improvement, rather than verifying $f=\sum^{N}f_{i}\lambda_{i}$, is to verify $0=(-1)f+\sum^{N}f_{i}\lambda_{i}$, so there is no output to construct. It would also be possible to run the heaps with minimal degree at the root (i.e. verifying the trailing coefficient first), which might be more advantageous in the case where there is actually an error, i.e. $0\neq(-1)f+\sum^{N}f_{i}\lambda_{i}$.
IV-D Interaction with geobuckets
[Joh74] assumed that the $g_{i}$ are stored as lists: what if they are stored as geobuckets? Then the process of replacing $g$ by $g-g_{1}$ is no longer $O(1)$, but depends on how the overall geobucket structure is stored: if it’s a vector of $b$ buckets, then the cost is $O(b)=O(\log\#g_{i})$. There are (at least) three possible routes.
  1. 1.
    Do an initial conversion of $g$ to list structure, adding the buckets from small to large, and then using [Joh74]. The cost is #$\#g$ in time (and space!).
  2. 2.
    Instead of storing $f_{i}g$ in the [Joh74] heap, store each $f_{i}g^{(j)}$ where the $g^{(j)}$ are the buckets of $g$. This means there are $b$ times as many entries in the heap, so multiplies the cost of [Joh74] by $O(\log b)=O(\log\log\#g)$
  3. 3.
    In practice we would probably do a hybrid: add the small buckets together but store $f_{i}g^{(j)}$ separately for large $g^{(j)}$, where “large” is a tuning parameter.